OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? x^4 -52x^2 + 576= 0 A. 4, -4 B. 4, -6 c. 4, -4, 6, -6 d. no solution

OpenStudy (helder_edwin):

u could factor the equation like this \[\large x^4-52x^2+576=(x^2-a)(x^2-b) \] such that \(a\cdot b=576\) and \(a+b=52\)

OpenStudy (anonymous):

Another way (maybe easier), would be to solve by substitution: Let \[u=x^{2}\] then \[x ^{4}-52x^2+576 = u^2-52u+576\] Then use the quadratic formula: \[u= \frac{ 52 \pm \sqrt{(-52)^2-4*1*576}}{ 2 }\] \[u=36, 16\] Substitute back in for u: \[u = x^2 = 36, 16\] \[x = \pm6, \pm4\]